A 3 gallon 20% alcohol water mixture is to be upgraded by adding to it a certain amount of 70% alcohol water mixture. How many gallons of the 70% must be added to make new enriched mixture of a 40% alcohol water mix?
Answer by mananth(16946)

--- percent ---------------- Amount
Alcool I 20 ---------------- 3 gal
Alcohol II 70 ---------------- x gal
Mixture 40 ---------------- 3 + x gal

20 * 3 + 70 x = 40 ( 3 + x )
60 + 70 x = 120 + 40 x
70 x -40 x = -60 + 120
30 x = 60
/ 30
x = 2 gal Alcohol II

m.ananth@hotmail.ca